1. How many grams of Lactic acid (C6H10O5) should be dissolved in 650g of cyclohexane (C6H12) to raise the boiling point to 84.910C? The normal boiling point of cyclohexane is 810C and its molal boiling point is 2.75 0C/m.




1
Expert's answer
2022-02-19T06:51:35-0500

!!! C6H10O5 is lactic acid lactate because lactic acid is C3H6O3 !!!


Solution:

Δt = i × Kb × m

where:

Δt = the change in boiling point

i = Vant Hoff's coefficient

Kb = the molal boiling-point elevation constant

m = the molality of the solute


Δt = 84.91°C - 81°C = 3.91°C

i = 1 (for lactic acid lactate - nonelectrolyte)

Kb = 2.75°C/m (for cyclohexane)


Therefore,

m = Δt / (i × Kb)

m = (3.91°C) / (1 × 2.75°C/m) = 1.422 m

Molality of C6H10O5 = 1.422 m


Molality of C6H10O5 = Moles of C6H10O5 / Kilograms of cyclohexane

Therefore,

Moles of C6H10O5 = Molality of C6H10O5 × Kilograms of cyclohexane

Moles of C6H10O5 = (1.422 m) × (0.65 kg) = 0.9243 mol


The molar mass of lactic acid lactate (C6H10O5) is 162.14 g/mol

Therefore,

Mass of C6H10O5 = (0.9243 mol) × (162.14 g / 1 mol) = 149.87 g

Mass of C6H10O5 = 149.87 g


Answer: 149.87 grams of C6H10O5 should be dissolved.